40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 25 02:03:59 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts 
    Reply  Fri Jul 28 15:36:32 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts errSig.pdfQPD_simulink.pngfeedbackTopology.pdf
       Reply  Tue Aug 1 16:05:01 2017, gautam, Update, Optical Levers, Optical lever tuning - cost function construction loopPerformance.pdfOplevLoop.pngOL_errSigs.pdfDAC_saturation.pdf
          Reply  Thu Nov 16 13:57:01 2017, gautam, Update, Optical Levers, Optical lever noise VertexOLnoise.pdfITMYvsSRM.pdf
             Reply  Thu Nov 16 15:43:01 2017, rana, Update, Optical Levers, Optical lever noise 
                Reply  Tue Nov 21 11:37:29 2017, gautam, Update, Optical Levers, BS OL calibration updated OL_calib_BS_PERROR.pdfOL_calib_BS_YERROR.pdfVertexOLnoise_updated.pdf
                   Reply  Tue Nov 21 16:28:23 2017, gautam, Update, Optical Levers, BS OL calibration updated BS_oplev_sweep.png
                      Reply  Tue Nov 21 23:04:12 2017, gautam, Update, Optical Levers, Oplev "noise budget" vertexOL_noises.pdf
                         Reply  Wed Nov 22 05:41:32 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                            Reply  Wed Nov 22 15:29:23 2017, gautam, Update, Optical Levers, Oplev "noise budget" OplevTest.jpg
                               Reply  Wed Nov 22 16:40:00 2017, Koji, Update, Optical Levers, Oplev "noise budget" 
                               Reply  Wed Nov 22 19:20:01 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                                  Reply  Wed Nov 22 23:56:14 2017, gautam, Update, Optical Levers, Oplev "noise budget" 
                                     Reply  Thu Nov 23 18:03:52 2017, gautam, Update, Optical Levers, Oplev "noise budget" ITMY_P_noise.pdfITMY_Y_noise.pdf
Message ID: 13439     Entry time: Tue Nov 21 16:28:23 2017     In reply to: 13437     Reply to this: 13441
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: BS OL calibration updated 

The numbers I have from the fitting don't agree very well with the OSEM readouts. Attachment #1 shows the Oplev pitch and yaw channels, and also the OSEM ones, while I swept the ASC_PIT offset. The output matrix is the "naive" one of (+1,+1,-1,-1). SUSPIT_IN1 reports ~30urad of motion, while SUSYAW_IN1 reports ~10urad of motion.

From the fits, the BS calibration factors were ~x8 for pitch and x12 for yaw - so according to the Oplev channels, the applied sweep was ~80urad in pitch, and ~7urad in yaw.

Seems like either (i) neither the Oplev channels nor the OSEMs are well diagonalized and that their calibration is off by a factor of ~3 or (ii) there is some significant imbalance in the actuator gains of the BS coils...

Quote:
 

Need to double check against OSEM readout during the sweep.

 

Attachment 1: BS_oplev_sweep.png  34 kB  | Show | Show all
ELOG V3.1.3-